OpenStudy (anonymous):

What are the solution(s) to the following system of equations? y = x2 – 1 y = x – 1 (0, 0) and (–1, 0) (0, –1) and (1, 0) (1, 0) and (0, 1) (1, 1) and (–1, –1)

OpenStudy (mandre):

x2 - 1 = y and x - 1 = y so you can set x2-1 = x-1 Solve that then substitute those answers into one of your first 2 equations to get your equivalent solutions for y.

OpenStudy (amistre64):

y = x2 – 1 y = x – 1 x2 – 1 = x – 1 x2 = x x2-x = 0 x(x-1) = 0

OpenStudy (anonymous):

or just plug each coordinate in and see which equal each other

OpenStudy (mandre):

@Sammy720 I would rather learn how to find answers than just try everything to see which one fits. You don't learn anything that way. Yes, I told someone else to do that but I won't be doing that again because it is not good advice.